OpenStudy (anonymous):

Solve for the variable. x/6=6.8

Parth (parthkohli):

Try to multiply both sides by 6.

Parth (parthkohli):

\[ \color{Black}{\Rightarrow {x \over 6} \times 6 = 6.8 \times 6 }\] Got it?

Parth (parthkohli):

Note that the 6's cancel out.

OpenStudy (anonymous):

need an answer!

OpenStudy (anonymous):

it is 40.8 rite??

Parth (parthkohli):

Can you calculate \(6.8 \times 6\)?